OpenStudy (anonymous):

The cost of Jaime's heating bill in January was 120% of the cost of her heating bill in February. Jaime's heating bill in January was $90. What is the cost of her heating bill in February?

OpenStudy (jdoe0001):

grab the correspondent components 100% in February corresponds to 120% in January $90 in Februarywill correspond to "some amount we dunno" in January in ratio terms \(\bf \cfrac{100}{120}=\cfrac{90}{x}\) solve for "x"

OpenStudy (anonymous):

90 times 120 is 10800. So 10,800 divided by 100 is 108.

OpenStudy (jdoe0001):

yes, you can see it this way, 20% is really 20/100 = 1/5 now what's 1/5 of 90? well, 1/5 of 90 => 90/5 = 18 so 18 is 20% of 90 90 + 18 = 108 :)

OpenStudy (jdoe0001):

using ratio terms, just put the corresponding ones in a fraction like above, and solve by the "unknown"

OpenStudy (anonymous):

Ohhhh thanks!

OpenStudy (jdoe0001):

yw